# Fast Ewald Summation using Prolate Spheroidal Wave Functions

### v1

@@ B.3 Evaluation of $\psi_{0}^{c}$ and $\lambda_{0}$
using high-order Gauss–Legendre quadrature [23, Sec. 3.2].
@@ B.5 Truncation order
The required number of terms $K$ in the expansion grows linearly with the bandlimit $c$. Empirically, $K\approx 1.2c$ suffices to achieve $10^{-12}$ accuracy, in agreement with the decay estimates for the Legendre coefficients in [23, Remark 9].
Appendix C Auxiliary standard results
C.1 Radial extension (Fourier-side definition)
@@ C.2 Derivation of (19)
C.3 A lattice-to-continuum estimate
We recall a standard lattice covering argument that bounds radial lattice sums by corresponding integrals. This estimate is used in the proof of Lemma 4 to pass from discrete sums to radial integrals.

### v2

@@ B.3 Evaluation of $\psi_{0}^{c}$ and $\lambda_{0}$
using high-order Gauss–Legendre quadrature [24, Sec. 3.2].
@@ B.5 Truncation order
The required number of terms $K$ in the expansion grows linearly with the bandlimit $c$. Empirically, $K\approx 1.2c$ suffices to achieve $10^{-12}$ accuracy, in agreement with the decay estimates for the Legendre coefficients in [24, Remark 9].
Appendix C Proofs of Lemmas 3 and 4
Proof of Lemma 3.
The Rayleigh quotient associated with (30) is
$\chi_{0}=\min_{\varphi\neq 0}\frac{\int_{-1}^{1}(1-s^{2})|\varphi^{\prime}(s)|^{2}\,\mathrm{d}s+c^{2}\int_{-1}^{1}s^{2}|\varphi(s)|^{2}\,\mathrm{d}s}{\int_{-1}^{1}|\varphi(s)|^{2}\,\mathrm{d}s}.$ (150)
Then, using the test function $\varphi(s)=1$,
$\chi_{0}\leq\frac{c^{2}\int_{-1}^{1}s^{2}\,\mathrm{d}s}{\int_{-1}^{1}1\,\mathrm{d}s}=\frac{c^{2}}{3}.$ (151)
Proof of Lemma 4.
Let $u(s):=s\psi_{0}^{c}(s)$. Substituting $\psi_{0}^{c}=u/s$ into (30), we obtain
$pu^{\prime\prime}+p^{\prime}u^{\prime}+qu=0,$ (152)
where
$p(s):=\frac{s^{2}-1}{s^{2}},\qquad q(s):=\frac{s^{2}(c^{2}s^{2}-\chi_{0})-2}{s^{4}}=c^{2}-\chi_{0}s^{-2}-2s^{-4}.$ (153)
For $s>1$, define the Sonin-type functional (see, e.g., [32, p. 166])
$F(s):=u(s)^{2}+\frac{p(s)}{q(s)}(u^{\prime}(s))^{2}.$ (154)
Since $p(s)>0$ for $s>1$, it remains to verify that $q(s)>0$ on $(1,\infty)$.
Using Lemma 3, $\chi_{0}\leq c^{2}/3$, and assuming $c\geq\sqrt{3}$, we obtain for $s>1$,
$q(s)=c^{2}-\chi_{0}s^{-2}-2s^{-4}>c^{2}-\chi_{0}-2\geq\tfrac{2}{3}c^{2}-2\geq 0.$ (155)
Hence, $q(s)>0$ for all $s>1$, and therefore $F$ is well defined and nonnegative on $(1,\infty)$.
Using (152), in particular $pu^{\prime\prime}=-p^{\prime}u^{\prime}-qu$, differentiation yields
$F^{\prime}(s)=-\frac{(pq)^{\prime}(s)}{q(s)^{2}}(u^{\prime}(s))^{2}.$ (156)
Moreover,
$(pq)^{\prime}(s)=\frac{2(c^{2}+\chi_{0})s^{4}-4(\chi_{0}-2)s^{2}-12}{s^{7}}.$ (157)
For $s>1$, the numerator is bounded below by
$2(c^{2}+\chi_{0})-4(\chi_{0}-2)-12=2(c^{2}-\chi_{0}-2)\geq 2\left(\frac{2}{3}c^{2}-2\right)\geq 0,$ (158)
where we again used Lemma 3 and $c\geq\sqrt{3}$. Hence, $(pq)^{\prime}(s)\geq 0$ for all $s>1$, and therefore $F^{\prime}(s)\leq 0$ on $(1,\infty)$.
Thus, $F$ is nonincreasing on $(1,\infty)$. Therefore,
$u(s)^{2}\leq F(s)\leq\lim_{t\to 1^{+}}F(t)=u(1)^{2}=\bigl(\psi_{0}^{c}(1)\bigr)^{2}.$ (159)
It follows that
$|\psi_{0}^{c}(s)|\leq\frac{\psi_{0}^{c}(1)}{s}.$ (160)
Appendix D Auxiliary standard results
D.1 Radial extension (Fourier-side definition)
@@ D.2 Derivation of (19)
D.3 A lattice-to-continuum estimate
We recall a standard lattice covering argument that bounds radial lattice sums by corresponding integrals. This estimate is used in the proof of Proposition 1 to pass from discrete sums to radial integrals.